2. (a) In lecture, we saw Euler’s Totient Function Φ(n), which is defined as the number of positive integers less than or equal to n that are relatively prime to n. Suppose I want to compute Φ(84), as per lecture. I know that 84 = 14 × 6, so I compute Φ(14) and multiply it by Φ(6). Do I get the right result? Briefly, why does this work or not work? Your answer should be brief, but not as simple as “because the numbers do/don’t multiply to form that.” (b) Compute Φ(84). Show your work.
Answer #1

(84) = 24
(6) = 2 {1,5}
(14) = 6   {1,3,5,9,11,13}

Reason -> Euler's totient function is a multiplicative function.It means if there are two numbers a and b are relatively prime then
(m*n) = (m)*(n).
And in above case
gcd(6,14) = 2 //greatest common divisor
it means 6 and 14 are not relatively prime.So here,
(84) is not equal to (14)*(6)

24 != 2*6

24 != 12

for more understanding lets take 84 = 21*4 {21,4 are relatively prime gcd(4,21) == 1}
(4) = 2{1,3}
(21) =12
(84) = (4)*(21)
24 = 2*12
24 = 24

If you have any doubt.Please feel free to ask.Thanks